length of a wire of a potentiometer is 100 cm and the emf is e voltemf of a battery whose internal resistance 0.5 ohm.if balanced point l=30.emf of the batterry is?
 

Grade:12th pass

1 Answers

Arun
5 years ago
By principle of potentiometerE1/E=L1/LSince null point is obtained at 30 cmL1=30L=100 because when E1 is not connected null point is obtained at 100 cmE1/E=30/100E1=30E/100